# A Data-Centric Perspective on Evaluating Machine Learning Models for Tabular Data

### v2

@@ Abstract
Tabular data is prevalent in real-world machine learning applications, and new models for supervised learning of tabular data are frequently proposed. Comparative studies assessing the performance of models typically consist of model-centric evaluation setups with overly standardized data preprocessing. This paper demonstrates that such model-centric evaluations are biased, as real-world modeling pipelines often require dataset-specific preprocessing and feature engineering. Therefore, we propose a data-centric evaluation framework. We select 10 relevant datasets from Kaggle competitions and implement expert-level preprocessing pipelines for each dataset. We conduct experiments with different preprocessing pipelines and hyperparameter optimization (HPO) regimes to quantify the impact of model selection, HPO, feature engineering, and test-time adaptation. Our main findings are: 1. After dataset-specific feature engineering, model rankings change considerably, performance differences decrease, and the importance of model selection reduces. 2. Recent models, despite their measurable progress, still significantly benefit from manual feature engineering. This holds true for both tree-based models and neural networks. 3. While tabular data is typically considered static, samples are often collected over time, and adapting to distribution shifts can be important even in supposedly static data. These insights suggest that research efforts should be directed toward a data-centric perspective, acknowledging that tabular data requires feature engineering and often exhibits temporal characteristics. Our framework is available under: https://github.com/atschalz/dc_tabeval.
@@ 2 Related Work
Limitations of current evaluation frameworks.    Several benchmarks exist for evaluating tabular machine learning models, focusing on general model comparisons [11, 31, 29, 35, 63] and specific sub-problems [46, 28, 21, 74, 27]. However, these benchmarks do not provide preprocessing settings for the included datasets. Consequently, most studies adopt a fixed, standardized preprocessing for all datasets to concentrate on model comparisons [75, 34, 63, 35, 47]. While this model-centric approach is suitable for AutoML, it limits the real-world transferability of model comparisons, as models in practical applications typically follow dataset-specific preprocessing and feature engineering [81, 87, 39]. Our evaluation framework is the first to explicitly incorporate a more detailed distinction through diverse preprocessing pipelines. Furthermore, existing benchmarks lack an external reference (e.g., a leaderboard) for the current best task performance, hindering comparability across different studies. In contrast, we leverage datasets from ML competitions as an external benchmark for high performance on tasks. Many existing evaluation frameworks prioritize usability at the expense of representativeness by limiting sample sizes and removing high-cardinality categorical features, thus evaluating models on artificially constrained dataset versions [11, 35]. Our evaluation framework solely consists of tasks meaningful to the real world without imposing artificial restrictions on datasets. Finally, most evaluation frameworks concentrate on tasks where samples are identically and independently distributed (i.i.d.). However, distribution shifts are prevalent in many machine learning applications [51, 84, 86, 58, 91, 28], and adapting to these shifts in tabular data has received limited attention [45, 28]. In this paper, we point out that excluding tabular data with temporal characteristics undermines the reliability of benchmarks, as many real-world applications using the benchmarked models include such data.
@@ 3 A Data-Centric Evaluation Framework for Tabular Machine Learning
We propose an evaluation framework built upon three crucial aspects that are often overlooked in tabular data research: 1) Evaluation on realistic datasets without removing frequently occurring challenging aspects like high cardinality categorical features, 2) Dataset-specific expert preprocessing pipelines, and 3) Evaluation against human expert performance on hidden test sets. Figure 1 depicts an overview of our framework. Our design choices are additionally justified by the fact that for each dataset, at least one model in our evaluation ranks among the top 1% of all competition participants.
@@ 3.1 Collection of Relevant and Challenging Datasets
Figure 2: Illustration of the dataset selection process. Details on the criteria and all screened datasets can be found in the Appendix. The Figure only lists the competitions as temporal, which were not already excluded for other reasons. Consistent with related work, we include competitions that have timestamps but can be approached without time-sensitive feature engineering. In total, we identified 46 competition datasets with temporal characteristics (i.e., timestamps as a feature).
Name Year N (Train) N (Test) D (Raw/FE) Categorical Metric Model TTA
MBGM [5] 2017 4,209 4,209 377 / 59 8 / 47 r2 XGBoost No
SVPC [60] 2018 12,296 49,342 4,992 / 1,420 0 / 0 rmsle LGBM No
AEAC [8] 2013 32,769 58,921 9 / 315 9 / 7,518 auc Ensemble Yes
OGPCC [9] 2015 61,878 144,368 93 / 104 0 / 0 logloss Ensemble Yes
SCS [78] 2016 76,020 75,818 370 / 224 0 / 0 auc Ensemble Yes
BPCCM [6] 2016 114,321 114,393 132 / 313 19 / 18,210 logloss XGBoost No
SCTP [64] 2019 200,000 200,000 200 / 600 0 / 0 auc NN Yes
HQC [23] 2015 260,753 173,836 299 / 300 29 / 868 auc XGBoost No
IFD [2] 2019 590,540 506,691 432 / 263 49 / 13,553 auc CatBoost Yes
PSSDP [1] 2017 595,212 892,816 57 / 53 8 / 104 gini NN Yes
Table 1: Datasets included in our framework. N denotes sample size, and D dimensionality of the raw data and after expert feature engineering. Categorical lists the no. of categorical features and the no. of clusters of the highest-cardinality categorical feature. Metric corresponds to the competition metric. MBGM and SVPC are regression tasks, OGPCC is a multi-class classification task, and the remaining are binary classification tasks. Model corresponds to the best single model used in the original expert solution. For some datasets, no best single model could be distinguished due to the heavy ensembling used. TTA denotes if test-time feature engineering has been used in the implemented expert solution.
@@ 3.2 Expert Solutions and Preprocessing Pipelines
Our proposed evaluation framework includes three preprocessing pipelines. One closely resembles the pipelines researchers currently use for model evaluation, and the other two are dataset-specific and directly derived from expert solutions. All preprocessing pipelines are model-agnostic. Model-specific preprocessing steps are considered part of the model and are explained in the Appendix.
@@
Expert Feature Engineering    We select one high-performance expert solution from Kaggle for each dataset. The solution was chosen based on the private leaderboard rank and the descriptions’ quality and sufficiency. For each solution, we separate the data preparation from the remaining parts of the solution. For most datasets, this pipeline solely consists of feature engineering techniques. Besides a few distinctions between tree-based and deep learning models, the pipelines are model-agnostic. Model-specific preprocessing steps are considered part of the model in our framework and are explained in the Appendix. This paper focuses on a pipeline perspective and does not discuss single feature engineering steps further. Implementation details and feature engineering techniques used for specific datasets are provided in the Appendix and in our publicly available code. For this pipeline, we ensured that all feature engineering operations included were on the training data and that a model could have learned the same patterns without external information.
Test-Time Adaptation    This pipeline is exactly the same as the expert feature engineering pipeline, with the key difference that the test data is used for feature engineering where applicable. Most ML competitions are organized so that the test features (but not the targets) are given. We found that the top solutions used the test data in their data preparation for six of the datasets in our framework. Hence, this pipeline represents the actual preprocessing used by the experts. While this might be considered an unfair and unrealistic setup, there are applications where using unlabeled test data for unsupervised learning is applicable (see Appendix). We argue that this conceptualization makes many tabular ML competitions a test-time adaptation (TTA) task. TTA is a type of domain adaptation where test samples are used at test time in an unsupervised or self-supervised way to update or retrain a model [89, 50, 49, 76, 67]. We term the common Kaggle practice of engineering domain-invariant features at test time as ’test-time feature engineering’, which can be considered a type of test-time training [80, 55, 66]. With this preprocessing pipeline, we are the first to closer examine test-time feature engineering in Kaggle competitions.
@@ 3.3 Modeling and Evaluation Framework
Modeling Pipeline and Models    We implement a unified modeling pipeline for all datasets with a dataset-specific cross-validation (CV) ensembling procedure. The validation sets are used for early stopping and determining the best hyperparameters. The final test data predictions are an ensemble of averaging the test predictions of each fold. Our experiments compare 7 models and one AutoML solution for all datasets and preprocessing pipelines. We use three gradient-boosted tree libraries (XGBoost [20], LightGBM [43], and CatBoost [70]) because each was used in at least one of the expert solutions. Each expert solution that used neural networks developed a highly customized network for the particular competition. We want to assess whether recently developed general-purpose architectures can replace the high effort of building custom networks. Hence, we chose ResNet and FTTransformer [34] because they have been frequently used in recent benchmark comparisons and have shown strong performance [35, 63]. Because the Resnet essentially is an MLP with skip connections, it serves as a baseline representing what was already possible before the recent developments in DL for tabular data. In addition, we use two more recent approaches: MLP-PLR [32], which can help learn high-frequency functions, mitigating a major weakness of deep learning for tabular data [35]; and GRANDE [61], a recent representative of hybrid neural-tree models. We are aware that even more recent architectures exist. However, our focus is not on particular models but on datasets and preprocessing. To assess how well fully automated solutions perform without any preprocessing, we additionally evaluate AutoGluon, which has been shown to be the current best AutoML solution [30].
@@
Evaluation    We use the Kaggle API to automatically submit predictions and retrieve performance results after evaluating against the hidden targets. Each dataset is evaluated on the metric specified by the competition host. Instead of reporting this metric directly, we report the solution’s private leaderboard position as the percentile. This has the benefit that although different metrics are used to evaluate the model, comparisons across datasets are possible. In the Appendix, we additionally report performances on the actual metrics for each dataset. Throughout the paper, higher values represent a better performance (leaderboard position).
Figure 3: Performance gains from different modeling components on the private Kaggle leaderboard by dataset and model. Higher values correspond to a better position. ’Default’ corresponds to the model performance with default hyperparameters in a standardized preprocessing pipeline. Light and extensive HPO correspond to tuning hyperparameters in the same preprocessing pipeline. Expert FE and FE-TTA correspond to the model performance with extensively tuned hyperparameters in the feature engineering and the test-time adaptation pipeline respectively.
@@ 4 Experimental Evaluation
Our framework allows us to assess the dataset-specific individual performance impact of model selection, hyperparameter optimization, feature engineering, and test-time adaptation. As a general overview, Figure 3 shows how each of the analyzed modeling components improves over the default baseline for each model and dataset. The results demonstrate the importance of an external performance reference: If we only considered the standardized evaluation setup (blue/orange/green bars), we would only be scratching the surface of achievable task performance for many data sets.
@@ 4.1 How Model Comparisons Change When Considering Dataset-specific Preprocessing
Figure 4: Average leaderboard position of models with different preprocessing. Black horizontal lines denote the Spearman correlation between all experiments with the respective preprocessing.
Through the implemented modeling pipelines, it becomes possible to evaluate how model comparisons change when evaluating inside the typically used standardized pipelines vs. expert pipelines with and without test-time adaptation. Three observations stand out when evaluating models in different preprocessing pipelines (Figure 4). 1) The model rankings change considerably, as indicated by the relatively low Spearman coefficients between the standardized preprocessing pipeline and the other pipelines. 2) The performance gap between all models diminishes when considering expert preprocessing. On average, all models benefit from feature engineering, and multiple models can reach top performance. While all models benefit from TTA, the performance increase varies. 3) The superiority of CatBoost vanishes when considering dataset-specific preprocessing. The reason is that CatBoost already incorporates specific feature engineering steps in its algorithm for which other models need manual engineering, as we will further elaborate in Subsection 4.3.
@@ 4.2 Measurable Progress Through Recent Efforts
Figure 5: Progress made through recent models, illustrated by retrospective comparison to the Kaggle leaderboard.
Figure 5 shows the model ranking on the private Kaggle leaderboard when trained after standardized preprocessing. CatBoost achieves top ranks in three competitions (MBGM, BPCCM, HQC) where a high manual effort in feature engineering was previously necessary. Similar to Erickson et al., [25], AutoGluon achieves top ranks in two of these (BPCCM, HQC) and one additional competition (OGPCC). Regarding neural networks, novel architectures outperform the ResNet baseline on nine datasets and even outperform tree-based solutions for three datasets (SVPC, SCS, SCTP). All neural networks originally used in the competitions were custom-designed for the particular competition. Hence, our analysis confirms that meaningful progress has been made in developing general-purpose architectures for tabular data. Although the progress in the tabular data field is clearly visible, top performance cannot be reached without human effort for six datasets.

### v3

@@ Abstract
Tabular data is prevalent in real-world machine learning applications, and new models for supervised learning of tabular data are frequently proposed. Comparative studies assessing the performance of models typically consist of model-centric evaluation setups with overly standardized data preprocessing. This paper demonstrates that such model-centric evaluations are biased, as real-world modeling pipelines often require dataset-specific preprocessing, which includes feature engineering. Therefore, we propose a data-centric evaluation framework. We select 10 relevant datasets from Kaggle competitions and implement expert-level preprocessing pipelines for each dataset. We conduct experiments with different preprocessing pipelines and hyperparameter optimization (HPO) regimes to quantify the impact of model selection, HPO, feature engineering, and test-time adaptation. Our main findings are: 1. After dataset-specific feature engineering, model rankings change considerably, performance differences decrease, and the importance of model selection reduces. 2. Recent models, despite their measurable progress, still significantly benefit from manual feature engineering. This holds true for both tree-based models and neural networks. 3. While tabular data is typically considered static, samples are often collected over time, and adapting to distribution shifts can be important even in supposedly static data. These insights suggest that research efforts should be directed toward a data-centric perspective, acknowledging that tabular data requires feature engineering and often exhibits temporal characteristics. Our framework is available under: https://github.com/atschalz/dc_tabeval.
@@ 2 Related Work
Limitations of current evaluation frameworks.    Several benchmarks exist for evaluating tabular machine learning models, focusing on general model comparisons [4, 22, 21, 26, 55] and specific sub-problems [40, 20, 13, 63, 18]. However, these benchmarks do not provide preprocessing settings for the included datasets. Consequently, most studies adopt a fixed, standardized preprocessing for all datasets to concentrate on model comparisons [64, 25, 55, 26, 41]. While this model-centric approach is suitable for AutoML, it limits the real-world transferability of model comparisons, as models in practical applications typically follow dataset-specific preprocessing pipelines containing feature engineering techniques [68, 74, 31]. Our evaluation framework is the first to explicitly incorporate a more detailed distinction through diverse preprocessing pipelines. Furthermore, existing benchmarks lack an external reference (i.e., a leaderboard) for the current best task performance, hindering comparability across different studies. In contrast, we leverage datasets from ML competitions as an external benchmark for high performance on tasks. Many existing evaluation frameworks prioritize usability at the expense of representativeness by limiting sample sizes and removing high-cardinality categorical features, thus evaluating models on artificially constrained dataset versions [4, 26]. Our evaluation framework solely consists of tasks meaningful to the real world without imposing artificial restrictions on datasets. Finally, most evaluation frameworks concentrate on tasks where samples are identically and independently distributed (i.i.d.). However, distribution shifts are prevalent in many machine learning applications [45, 71, 73, 51, 78, 20], and adapting to these shifts in tabular data has received limited attention [39, 20]. In this paper, we point out that excluding tabular data with temporal characteristics undermines the reliability of benchmarks, as many real-world applications using the benchmarked models include such data.
@@ 3 A Data-Centric Evaluation Framework for Tabular Machine Learning
We propose an evaluation framework built upon three crucial aspects that are often overlooked in tabular data research: 1) Evaluation on realistic datasets without removing frequently occurring challenging aspects like high cardinality categorical features, 2) Dataset-specific expert preprocessing pipelines containing feature engineering techniques, and 3) Evaluation against human expert performance on hidden test sets. Figure 1 depicts an overview of our framework. Our design choices are additionally justified by the fact that for each dataset, at least one model in our evaluation ranks among the top 1% of all competition participants (note that not all participants are experts, and leaderboard distributions can vary across datasets).
@@ 3.1 Collection of Relevant and Challenging Datasets
Figure 2: Illustration of the dataset selection process. Details on the criteria and all screened datasets can be found in the Appendix. The Figure only lists the competitions as temporal, which were not already excluded for other reasons. In total, we identified 46 competition datasets with temporal characteristics (i.e., timestamps as a feature). Consistent with related work, we include competitions that have timestamps but can be approached without time-sensitive feature engineering.
Name Year N (train/test) D (raw/fe) Categorical Task Metric Model TTA
MBGM 2017 4209 / 4209 377 / 59 8 / 47 Reg r2 XGBoost No
SVPC 2018 12296 / 49342 4992 / 1420 0 / 0 Reg rmsle LGBM No
AEAC 2013 32769 / 58921 9 / 315 9 / 7,518 Bin auc Ensemble Yes
OGPCC 2015 61878 / 144368 93 / 104 0 / 0 Multi logloss Ensemble Yes
SCS 2016 76020 / 75818 370 / 224 0 / 0 Bin auc Ensemble Yes
BPCCM 2016 114321 / 114393 132 / 313 19 / 18,210 Bin logloss XGBoost No
SCTP 2019 200000 / 200000 200 / 600 0 / 0 Reg auc NN Yes
HQC 2015 260753 / 173836 299 / 300 29 / 868 Bin auc XGBoost No
IFD 2019 590540 / 506691 432 / 263 49 / 13,553 Bin auc CatBoost Yes
PSSDP 2017 595212 / 892816 57 / 53 8 / 104 Bin gini NN Yes
Table 1: Datasets included in our framework. N denotes the sample size in thousands, and D dimensionality of the raw data and after expert feature engineering. Categorical lists the no. of categorical features and the no. of clusters of the highest-cardinality categorical feature. Metric corresponds to the competition metric. MBGM and SVPC are regression tasks, OGPCC is a multi-class classification task, and the remaining are binary classification tasks. Model corresponds to the best single model used in the original expert solution. For some datasets, no best single model could be distinguished due to the heavy ensembling used. TTA denotes if test-time feature engineering has been used in the implemented expert solution.
@@ 3.2 Expert Solutions and Preprocessing Pipelines
In the context of our paper, preprocessing refers to a pipeline that combines a “set of techniques used prior to the application of a [model]” [19]. Feature engineering (FE) refers to techniques that “construct novel features from given data with the goal of improving predictive learning performance” [38]. Consequently, feature engineering is a subset of preprocessing. Our proposed evaluation framework includes three preprocessing pipelines. One is dataset-agnostic and closely resembles the pipelines researchers currently use for model evaluation. The other two are dataset-specific and directly derived from expert solutions. All preprocessing pipelines are model-agnostic, and model-specific preprocessing steps are considered part of the model in our framework.
@@
Expert Feature Engineering    We select one high-performance expert solution from Kaggle for each dataset. The solution was chosen based on the private leaderboard rank and the descriptions’ quality and sufficiency. For each solution, we separate the data preparation from the remaining parts of the solution. For most datasets, this pipeline solely consists of feature engineering techniques. Besides a few distinctions between tree-based and deep learning models, the pipelines are model-agnostic. Model-specific preprocessing steps (i.e., feature normalization for neural networks) are considered part of the model in our framework and are explained in the Appendix. This paper focuses on a pipeline perspective and does not discuss single feature engineering steps further. Implementation details and feature engineering techniques used for specific datasets are provided in the Appendix and in our publicly available code. For this pipeline, we ensured that all feature engineering operations included were on the training data and that a model could have learned the same patterns without external information. Some of the feature engineering techniques used by the experts occur across multiple datasets: groupby interactions of categorical and numeric features (4), 2-order categorical interactions (3), feature selection (3), categorical frequency encoding (3), dimensionality reduction (2), 3-order categorical interaction (2), 2-order arithmetic interactions (2), sum of missing values in a row (2), and sum of zeros in a row (2). A common pattern is that the most frequently applied feature engineering steps include categorical features and that feature interactions are frequently manually engineered while transformations of single features are rare.
Test-Time Adaptation    This pipeline is exactly the same as the expert feature engineering pipeline, with the key difference that the test data is used for feature engineering where applicable. Most ML competitions are organized so that the test features (but not the targets) are given. We found that the top solutions used the test data in their data preparation for six of the datasets in our framework. Hence, this pipeline represents the actual preprocessing used by the experts. While this might be considered an unfair and unrealistic setup, there are applications where using unlabeled test data for unsupervised learning is applicable (see Appendix A.3). We argue that this conceptualization makes many tabular ML competitions a test-time adaptation (TTA) task. TTA is a type of domain adaptation where test samples are used at test time in an unsupervised or self-supervised way to update or retrain a model [76, 44, 43, 65, 57]. We term the common Kaggle practice of engineering domain-invariant features at test time as test-time feature engineering. The feature engineering techniques most frequently used to this end are groupby interactions, frequency encodings, and learning joint low-dimensional representations. With this preprocessing pipeline, we are the first to closer examine test-time feature engineering in Kaggle competitions.
@@ 3.3 Modeling and Evaluation Framework
Modeling Pipeline and Models    We implement a unified modeling pipeline for all datasets with a dataset-specific cross-validation (CV) ensembling procedure. The validation sets are used for early stopping and determining the best hyperparameters. The final test data predictions are an ensemble of averaging the test predictions of each fold. We use three gradient-boosted tree libraries (XGBoost [12], LightGBM [36], and CatBoost [60]) because each was used in at least one of the expert solutions. Each expert solution that used neural networks developed a highly customized network for the particular competition. We want to assess whether recently developed general-purpose architectures can replace the high effort of building custom networks. Hence, we chose ResNet and FTTransformer [25] because they have been frequently used in recent benchmark comparisons and have shown strong performance [26, 55]. Because the Resnet essentially is an MLP with skip connections, it serves as a baseline representing what was already possible before the recent developments in DL for tabular data. In addition, we use two more recent approaches: MLP-PLR [23], which can help learn high-frequency functions, mitigating a major weakness of deep learning for tabular data [26]; and GRANDE [53], a recent representative of hybrid neural-tree models. Although other recent architectures exist, we don’t include more, as our focus is not on benchmarking particular models but rather on demonstrating the importance of data-centric evaluation. To assess how well fully automated solutions perform without any preprocessing, we additionally evaluate AutoGluon, which has been shown to be the current best AutoML solution [21].
@@
Evaluation    We use the Kaggle API to automatically submit predictions and retrieve performance results after evaluating against the hidden targets. Each dataset is evaluated on the metric specified by the competition host. Instead of reporting this metric directly, we report the solution’s private leaderboard position as the percentile. This has the benefit that although different metrics are used to evaluate the model, comparisons across datasets are possible. Note that the leaderboard position is always a snapshot of the end of each competition. In the Appendix, we additionally report performances on the actual metrics for each dataset. Throughout the paper, higher values represent a better performance (leaderboard position). As we use only one test set per dataset and less datasets compared to academic benchmarks, concerns about overfitting might be raised. However, ten datasets with one test set are less of an issue in our framework than it would be in conventional benchmarks, because: 1) The datasets in our framework, especially the test data, are comparably large and overfitting them is harder. Roelofs et al. [61] found that at least 10,000 test examples is a reasonable minimum test set size to protect against adaptive overfitting in Kaggle challenges. All test sizes in our framework, except for the MBGM dataset, are at least  of size 50,000. 2) Test labels are unknown making it hard to purposefully overfit on particular samples. 3) The need of submitting to Kaggle, although automated, is an additional overfitting barrier.
Figure 3: Performance gains from different modeling components on the private Kaggle leaderboard by dataset and model. Higher values correspond to a better position. Each segment represents the performance gain of adding the modeling component to the previous configuration. ’Default’ corresponds to the model performance with default hyperparameters in a standardized preprocessing pipeline. Light and extensive HPO correspond to tuning hyperparameters in the same preprocessing pipeline. Expert FE and FE-TTA correspond to the model performance with extensively tuned hyperparameters in the feature engineering and the test-time adaptation pipeline respectively.
@@ 4 Experimental Evaluation
Our framework allows us to assess the dataset-specific individual performance impact of model selection, hyperparameter optimization, feature engineering, and test-time adaptation. Whenever not stated otherwise, we report the results for the extensively tuned hyperparameter setup. As a general overview, Figure 3 shows how each of the analyzed modeling components improves over the default baseline for each model and dataset. The length of each segment indicates the performance gain relative to the previous configuration. The results demonstrate the importance of an external performance reference: If we only considered the standardized evaluation setup (blue/orange/green segments), we would only be scratching the surface of achievable task performance for many data sets.
@@ 4.1 How Model Comparisons Change When Considering Dataset-specific Preprocessing
Figure 4: Average leaderboard position of models with different preprocessing. Black horizontal lines denote the Spearman correlation between all results with the respective preprocessing.
Three observations stand out when evaluating models in different preprocessing pipelines (Figure 4). 1) The model rankings change considerably, as indicated by the relatively low Spearman coefficients between the standardized preprocessing pipeline and the other pipelines. 2) The performance gap between all models diminishes when considering expert preprocessing. On average, all models benefit from feature engineering, and multiple models can reach top performance. While all models benefit from TTA, the performance increase varies. 3) The superiority of CatBoost vanishes when considering dataset-specific preprocessing. The reason is that CatBoost already incorporates specific feature engineering steps in its algorithm for which other models need manual engineering, as we will further elaborate in Subsection 4.3.
@@ 4.2 Measurable Progress Through Recent Efforts
Figure 5: Progress made through recent models trained in the standardized preprocessing pipeline, illustrated by retrospective comparison to the Kaggle leaderboard. Best NN denotes the best model of FTTransformer, MLP-PLR, and GRANDE.
Figure 5 shows the model ranking on the private Kaggle leaderboard when trained after standardized preprocessing. CatBoost achieves top ranks in three competitions (MBGM, BPCCM, HQC) where a high manual effort in feature engineering was previously necessary. Similar to Erickson et al., [16], AutoGluon achieves top ranks in two of these (BPCCM, HQC) and one additional competition (OGPCC). Regarding neural networks, novel architectures rank higher than the ResNet baseline on nine datasets. In two competitions, neural networks were originally the single-best models (SCTP, PSSDP - see Table 1). In our analysis, MLP-PLR and FTTransformer are able to reach top ranks on these datasets after feature engineering and test-time adaptation, while ResNet performs worse (see Figure 3). All neural networks originally used in the competitions were custom-designed for the particular competition. Hence, our analysis confirms that meaningful progress has been made in developing general-purpose architectures for tabular data as they reduce the necessity of custom-designed networks. Although the progress in the tabular data field is clearly visible, top performance cannot be reached without human effort for six datasets.